# Compton Form Factor Extraction using Quantum Deep Neural Networks

### v2

@@ III.2.2 Quantum Qualifier
[…]
@@
In Tables 12, 13, and 14 in the Appendix, we show the values of these five function characteristics for each of the target functions with $0.1\sigma$, $0.25\sigma$, $1\sigma$ noise level, respectively. We then use Eq. (46) to calculate the quantum outperformance $\Xi$ for each of the tests, and we analyze the values of $\Xi$ for a given number of epochs against each of the five function characteristics. We perform linear regressions on this and display the results in Tables 15 and 16, which indicate the correlation between quantum outperformance and the characteristics of the target function data for a given number of epochs. Using this information, we fit polynomial curves (and an exponential curve for $\mathfrak{N}$) to the regression slopes and $R^{2}$ values as a function of the number of epochs. Since the polynomial fit of the regression slopes against the number of epochs approximates the correlation between quantum outperformance $\Xi$ and a given function characteristic, we construct a linear function corresponding to this correlation for a given number of epochs. Finally, we define the quantum qualifier $\hat{\Xi}$ by summing all of these linear functions (with slope given by the $\Xi$ regression slope polynomial fit) and then weighting by the strength of the correlation (given by the $R^{2}$ polynomial fit):
$\hat{\Xi}=e^{-\alpha n}\sum_{i=0}^{2}\beta_{1i}\,n^{i}X_{1}+\sum_{j=2}^{5}\sum_{i=0}^{4}\beta_{ji}n^{i}X_{j}.$ (55)
This quantum qualifier can estimate the value of $\Xi$ based on the five function characteristics and the number of epochs $n$ that the CDNN and QDNN are trained on. Here, the coefficients are parameterized in one indexed set $\beta_{ji}$, where $j$ indexes the corresponding observable $X_{j}$:
$X_{j}\in\{\mathfrak{N}-0.25,\;\Phi-24.5,\;\mathfrak{D}-0.95,\;\mathfrak{M}+0.05,\;\mathfrak{F}-4999.5\}.$ (56)
The exponential prefactor $e^{-\alpha n}$ is included explicitly (with $\alpha=0.0101$), and parameters $\beta_{ji}$ encode all coefficients. $\alpha$ is a fixed scalar in the exponential. The outer sum $j$ runs over different offset parameters ($\mathfrak{N},\Phi,\mathfrak{D},\mathfrak{M},\mathfrak{F}$), each scaled by a specific offset. The inner sum $\sum_{i}$ runs over powers of $n$, representing polynomial dependence. Table 17 in the Appendix shows the numerical parameters of the quantum qualifier expression.
Equation (55) captures the main conclusions that are immediately evident from Tables 15 and 16; for example, that mutual information $\mathfrak{M}$ has the most positive correlation with quantum outperformance, while fractal dimension $\mathfrak{D}$ has the most negative and strongest correlation with quantum outperformance. We will return to the idea of a quantum qualifier in Sec. VI, where we construct a simpler version specifically for the DVCS process, which we will use to help determine if a QDNN is advantageous in CFF extraction. For now, having thoroughly discussed the structure and application of QDNNs, we will detail the preparation of the data essential for training and validating these models in the context of CFF extraction. It is worth noting that this process of testing data and constructing a quantum qualifier can be done for any formalism, at any twist approximation, or for any set of observables, with sensitivity to error and data sparsity. The construction of such metrics may have great utility for other hadronic physics applications.
IV Experimental Data and Pseudodata Generation
The CFF extraction analyses detailed in the remainder of this paper utilize DVCS experimental data from Hall A [55, 54, 53] and Hall B [56] at Jefferson Lab (JLab). The dataset, summarized in Table 2, includes both helicity-independent and helicity-dependent cross sections. However, our focus is exclusively on the helicity-independent (unpolarized) cross sections for simplicity. The data is finely binned in a fourfold differential cross section parameterized by $Q^{2}$, $x_{B}$, $t$, and $\phi$, comprising a total of 3,885 data points for analysis.
@@ IV Experimental Data and Pseudodata Generation
IV.1 Pseudodata Generation
We now generate numerous pseudodata sets based directly on the experimental data sets. These pseudodata sets allow us to precisely control the “true” CFF values, offering a controlled environment for testing and evaluating the accuracy and performance of our neural network models before their application to real experimental data, where the true values are unknown. Our pseudodata sets are generated using the following basic model function:
@@ IV.1 Pseudodata Generation
where $G(x_{B},t)$ represents the CFFs and DVCS cross section components: $\Re e\mathcal{H},\Re e\mathcal{E},\Re e\widetilde{\mathcal{H}},$ and $DVCS$. The parameters $\{a,b,c,d,e,f\}$ for each case are listed in Table 3, and are based on initial tests. The structure of this model function allows for linear or non-linear changes in $x_{B}$ with exponential flexibility in $t$. The parameterization has no physical meaning as the model is purely designed for testing the DNN extractions.
@@
$\Re e\widetilde{\mathcal{H}}$1.86 1.50 -0.29 -1.33 0.46 -0.98
$DVCS$ 0.50 -0.41 0.05 -0.25 0.55 0.166
@@ V Pseudodata CFF Extraction
We compare the performance of CDNNs and QDNNs in extracting CFFs, assessing their effectiveness through a testing phase using pseudodata generated at the same kinematics as the experimental data. The main evaluation metrics we use are accuracy and precision. Accuracy $\epsilon(k,Q^{2},x_{B},t)$ quantifies how closely the mean of the DNN-extracted CFFs aligns with the true CFF values used in pseudodata generation:
@@
The generating function used to produce the true CFF values follows a basic model from Eq. (57) using parameters from Table 3, and the pseudodata we extract the CFFs from is described in Sec. IV.1.
In this section, we aim to compare CDNN and QDNN approaches, evaluating their ability to extract CFFs with improved accuracy, precision, and reliability. We consider two QDNN models: the Basic QDNN described in the next section (Sec. V.1) and the Full QDNN described in following section (Sec. V.2). The Basic QDNN is designed to mirror the CDNN model in architecture and complexity, serving as a baseline comparison without quantum-specific optimization. The Full QDNN is designed to fully take advantage of the quantum potential by carefully mitigating quantum-specific challenges and incorporating quantum-specific optimization schemes. The details of the architecture and construction of the CDNN and QDNNs, the method of extraction, and the corresponding results are presented in the following subsections.
@@ V.1 Basic QDNN
All the DNN models used for extraction in this paper have a three-dimensional input ($x_{B}$, $Q^{2}$, $t$), the kinematic values of a given kinematic bin, and a four-dimensional output ($\Re e\mathcal{H}$, $\Re e\mathcal{E}$, $\Re e\widetilde{\mathcal{H}}$, $DVCS$), the four CFFs. The DNNs use a custom, physics-informed loss function that predicts the cross section $F$ using the outputted three CFFs and DVCS term with the BKM10 formalism (Sec. II) and computes the mean squared error of the predicted cross section against the true cross section. The specific model architectures of the CDNN and QDNN are described as follows.
The CDNN model used for CFF extraction is based closely on the regression CDNN model described in Sec. III.2 and shown in Fig. 5. The input layer of the CDNN maps the three-dimensional input layer ($x_{B}$, $Q^{2}$, $t$) to 64 neurons. Then, there are 8 hidden layers, each with 64 neurons and ReLU activation. Finally, the output layer maps the 64 features to a four-dimensional output ($\Re e\mathcal{H}$, $\Re e\mathcal{E}$, $\Re e\widetilde{\mathcal{H}}$, $DVCS$). The QDNN is constructed based on the architecture of the CDNN and is similar to the architecture shown in Fig. 6. The input layer of the QDNN is a classical linear preprocessing layer, mapping the three-dimensional input to 6 qubits, which is equivalent to the 64 neurons used in the CDNN (Sec. III). The input data is then embedded into quantum state space using angle embedding, which encodes the features as rotation angles on the qubits. Next, there are 8 hidden Strongly Entangling Layers, consisting of rotations and entanglers. Finally, the expectation value of the Pauli-Z operator for each qubit is measured. In the classical postprocessing layers, the 6 quantum outputs are mapped to 64 neurons, followed by a Tanh activation, then another linear layer to the final four-dimensional output.
In the remainder of this section, we first perform a thorough analysis of the model complexities of the CDNN and QDNN in order to show quantitatively that the two models are similar in complexity (Sec. V.1.1). Then, we use the models to extract the CFFs from the pseudodata and perform a detailed error analysis (Sec. V.1.2).
@@ V.1.1 Model Complexity
To compare the CDNN and QDNN fairly, we must ensure that neither model has an inherent advantage simply due to size or computational power. Therefore, in this section, we compute values of several quantitative metrics that aim to characterize and compare the complexity of the CDNN and QDNN architectures employed for the extraction of CFFs from noisy pseudodata. Specifically, we focus on the number of trainable parameters and the estimated number of floating point operations per second (FLOPs), which serve as metrics for a model’s capacity and computational cost, respectively. The goal of this analysis is to demonstrate that the CDNN and Basic QDNN are comparable in the sense that they operating at roughly the same level of complexity in terms of the size and depth of the model, even though they come from very different paradigms.
However, even though we model the architecture of the Basic QDNN on that of the CDNN, making such a head-to-head comparison is not so straightforward because CDNNs and QDNNs are fundamentally different. For example, in a QDNN, the number of trainable parameters (e.g., rotation angles) may be relatively small, but the entangled structure of quantum circuits can, in principle, encode highly non-linear transformations that would require a much larger classical network to approximate. Similarly, estimating FLOPs in quantum circuits is nontrivial because quantum gates do not execute classical floating point operations, and instead evolve qubit states via unitary transformations, which scale differently depending on circuit depth and entanglement structure. Still, calculating these architectural metrics is a valuable starting point. They give us a baseline for comparing the CDNN and QDNN models side by side and ensure that we are not giving one model an unfair advantage simply by making it larger or more powerful.
For a fully connected layer with input size $M$ and output size $N$, the number of trainable parameters $P$ is
$P=MN+N\quad(\text{weights + biases}).$ (60)
For the CDNN, we have trainable parameters in the input layer (364)3$\to$ 64)3 → 64 ), 8 hidden layers (6464)64$\to$ 64)64 → 64 ), and output layers (644)64$\to$ 4)64 → 4 ). Therefore, the total number of trainable parameters in the CDNN is
$\begin{split}P_{\text{CDNN}}&=P_{\text{in}}+8P_{\text{hid}}+P_{\text{out}}\\ &=(3\cdot 64+64)+8(64\cdot 64+64)+(64\cdot 4+4)\\ &=33796.\end{split}$ (61)
For the QDNN, we have trainable parameters in the preprocessing layer (363$\to$ 63 → 6), 8 Strongly Entangling Layers ($\to 6$ 6), and the postprocessing layers (6644)6$\to$ 64$\to$ 4)6 → 64 → 4 ). The angle embedding layer is a fixed operation that contains no trainable parameters. In the Strongly Entangling Layers, each qubit undergoes three parameterized rotations, and since there are 8 Strongly Entangling Layers and 6 qubits per layer,
@@
Therefore,
$\begin{split}P_{\text{QDNN}}&=P_{\text{pre}}+P_{\text{SEL}}+P_{\text{post}}\\ &=(3\cdot 6+6)+144+((6\cdot 64+64)+(64\cdot 4+4))\\ &=876.\end{split}$ (63)
As expected, the number of trainable parameters for the QDNN is much lower than that of the CDNN because of its ability to encode a $2^{n}$-dimensional Hilbert space into $n$ qubits. A more direct comparison of CDNN and QDNN model complexity is FLOPs.
Floating Point Operations per Second (FLOPs) measure the computational cost of the forward pass through a DNN. For a single dense linear layer with input size $M$ and output size $N$, the number of FLOPs $C$ is
@@
For the CDNN, we have the input layer (3643$\to$ 643 → 64), a ReLU activation (64 FLOPs), 8 hidden layers ($\to 64$ 64) and associated ReLU activations (64 FLOPs), and an output layer (64464$\to$ 464 → 4). Therefore,
$\begin{split}C_{\text{CDNN}}&=(C_{\text{in}}+C_{\text{ReLU}})+8(C_{\text{hid}}+C_{\text{ReLU}})+C_{\text{out}}\\ &=(2\cdot 3\cdot 64+64)+8(2\cdot 64\cdot 64+64)+2\cdot 64\cdot 4\\ &=67008.\end{split}$ (65)
[…]
$\begin{split}C_{\text{QDNN}}&=C_{\text{pre}}+8C_{\text{SEL}}+(C_{\text{post}}+C_{\text{Tanh}})\\ &=(2\cdot 3\cdot 64)+8(2\cdot 64\cdot 64)\\ &\mathrel{\phantom{=}}+\,(2\cdot 64\cdot 64+64+2\cdot 64\cdot 4)\\ &=74688,\end{split}$ (66)
which is comparable to the CDNN.
The model complexity comparison results are summarized in Table 4. These results demonstrate that the CDNN and QDNN are similar in complexity and architecture and therefore serve as a good baseline comparison of the classical and quantum approaches to CFF extraction.
Figure 7: Distributions of the extracted $\Re e\mathcal{E}$ from noisy replicas of cross section pseudodata generated from kinematics $k=5.75$ GeV, $Q^{2}=2.22$ GeV2, $x_{B}=0.333$, $t=-0.16$ GeV2 (Set 144). The histograms display improved accuracy and precision by the QDNN.
(a) CDNN
(b) QDNN
Figure 8: Predicted cross sections from the CDNN and QDNN fits of noisy replicas of cross section pseudodata (red points) generated from kinematics $k=8.521$ GeV, $Q^{2}=3.65$ GeV2, $x_{B}=0.367$, $t=-0.20459$ GeV2 (Set 26).
(a) CDNN
(b) QDNN
Figure 9: Predicted cross sections from the CDNN and QDNN fits of noisy replicas of cross section pseudodata (red points) generated from kinematics $k=5.75$ GeV, $Q^{2}=2.48$ GeV2, $x_{B}=0.399$, $t=-0.45$ GeV2 (Set 165). In both cases there is a small improvement in proximity to the true with a clear narrowing of the error band for the QDNN.
@@ V.1.2 CFF Extraction Test
To quantify overall performance, we define a reduced $\chi^{2}$ metric, $M_{\chi^{2}}$, as the sum of the reduced $\chi^{2}$ values for each CFF across all 195 kinematic bins. A lower $M_{\chi^{2}}$ indicates a better fit, with a value of 4 corresponding to a perfect match. Averaging over all bins, we find $\overline{M}_{\chi^{2}}^{\text{CDNN}}\approx 78.97$ and $\overline{M}_{\chi^{2}}^{\text{QDNN}}\approx 28.18$, indicating that the QDNN outperforms the CDNN by approximately 64.3%, even without leveraging quantum-specific enhancements.
@@
The predicted cross sections using the predicted CFFs from the CDNN and QDNN are also analyzed. Using the framework described in Section II, we compute the predicted cross sections using the extracted CFFs and compare them to the true cross section values used to generate the pseudodata. In Figs. 8 and 9, we present results for two representative kinematic bins, plotting the mean predicted cross-section values as a function of the azimuthal angle $\phi$, along with their 68% confidence intervals. The true cross-section curve and fluctuated true cross-section points with simulated experimental error from the pseudodata sets are overlaid for reference. A key observation from these figures is that the improved precision of the QDNN extraction leads to narrower confidence intervals in the predicted cross sections across all three bins due to the reduced uncertainty in the extracted CFFs. In these examples, there is also a small, observable improvement in the accuracy of the predicted cross section by the QDNN.
@@
Understanding the sources of error is essential for robust model evaluation. A detailed analysis of the errors in the CFF extractions is necessary to fully understand and quantify the differences between QDNN and CDNN approaches. We consider four distinct error metrics: algorithmic error, methodological error, precision, and accuracy. To compute the algorithmic error, we perform the extraction on 1000 identical replicas of each bin and calculate the resulting spread of each CFF. Thus, algorithmic error represents error resulting from the limitations and inherent uncertainties of the extraction algorithm itself, rather than simulated experimental error in the data. To compute the methodological error, we introduce a spread in the parameters used to generate the CFFs (Table 3) and use the generating function [Eq. (57)] to generate CFFs sampled in a domain that approximately corresponds to the spread in the original extracted CFFs. Then, the methodological error is computed as the spread of the resulting residuals from the new extractions. Methodological error represents the uncertainty arising from the choice of model parameters and assumptions used in the extraction process. By introducing a spread in the parameters used to generate the CFFs and propagating this spread through the generating function, we assess how variations in these underlying assumptions influence the extracted CFF values. This approach allows us to estimate the systematic impact of modeling choices on the final extraction results, distinguishing it from purely statistical or algorithmic uncertainties. Finally, precision and accuracy are computed using Eqs. (59) and (58), respectively. These metrics provide an evaluation of the performance of the DNNs on the original pseudodata set, with precision measuring the consistency of the extracted values across multiple replicas and accuracy reflecting the agreement between the extracted and true CFFs. For all four of these error metrics, a lower value indicates a lower error.
@@
Figure 10 presents a comparative analysis of algorithmic error, methodological error, precision, and accuracy for the DNN-based extraction of each CFF across all 195 kinematic bins. In the plot, the absolute mean deviation from the true value is shown for each contribution as a point, while the spread in that deviation is represented by the error bar for each point. A notable improvement in the three CFFs’ precision is observed with the QDNN, whereas the other error contributions remain largely comparable between the two models. The QDNN shows an improved extraction accuracy for $\Re e\mathcal{E}$ and $\Re e\widetilde{\mathcal{H}}$ while the CDNN performs better for $\Re e\mathcal{H}$ and $DVCS$. Methodological errors are fairly small in both cases, but with a narrower distribution for the QDNN, indicating a bit better robustness to variations in the model parameters used to generate the CFFs. The increase in precision from the QDNN is also evident in the predicted cross-section plots (Figs. 8 and 9) as visibly tighter confidence intervals. Overall, QDNN demonstrates improved total error performance compared to CDNN, underscoring its effectiveness in this application.

### v3

@@ IV Experimental Data and Pseudodata Generation
Experiment Publication Year $E_{beam}$ (GeV) $Q^{2}$ (GeV2) t-t (GeV2) $x_{B}$ Number of Points
@@
The CFF extraction analyses detailed in the remainder of this paper utilize DVCS experimental data from Hall A [39, 40, 51] and Hall B [73] at Jefferson Lab (JLab). The dataset, summarized in Table 1, includes both helicity-independent and helicity-dependent cross sections. However, our focus is exclusively on the helicity-independent (unpolarized) cross sections for more direct comparison to previous work [30]. The data is finely binned in a fourfold differential cross section parameterized by $Q^{2}$, $x_{B}$, $t$, and $\phi$, comprising a total of 3,885 data points for analysis.
@@
We generate numerous pseudodata sets based directly on the experimental data sets. These pseudodata sets allow us to precisely control the “true” CFF values, offering a controlled environment for closure testing and evaluating the accuracy and performance of our neural network models before their application to real experimental data, where the true values are unknown. Our pseudodata sets are generated using the following basic model function:
@@
where $G(x_{B},t)$ represents the CFFs and DVCS cross section components: $\Re e\mathcal{H},\Re e\mathcal{E},\Re e\widetilde{\mathcal{H}}$, and $DVCS$. The parameters $\{a,b,c,d,e,f\}$ for each case are listed in Table 2, and are based on initial fits to experimental data. The structure of this model function allows for linear or non-linear changes in $x_{B}$ with exponential flexibility in $t$. The parameterization has no physical meaning as the model is purely designed for testing the DNN extractions independently of any theoretical model, so as to reduce the likelihood of bias in the testing phase.
@@
$\Re e\widetilde{\mathcal{H}} -$1.86 1.50 -0.29 -1.33 0.46 -0.98
DVCSDVCS 0.50 -0.41 0.05 -0.25 0.55 0.166
@@ V Pseudodata Comparison
We compare the performance of CDNNs and QDNNs in extracting CFFs, assessing their effectiveness through a testing phase using the pseudodata generated previously described. The main evaluation metrics used are accuracy and precision. Accuracy $\epsilon(k,Q^{2},x_{B},t)$ quantifies how closely the mean of the DNN-extracted CFFs aligns with the true CFF values used in pseudodata generation:
@@
The generating function used to produce the true CFF values follows a basic model from Eq. (19) using parameters from Table 2, and the pseudodata we extract the CFFs from is described in Sec. IV.
In the next subsections, we rigorously compare CDNN and QDNN approaches, evaluating their ability to extract CFFs with through accuracy, precision, and stability. Further, we consider two QDNN models: the Basic QDNN described in the next section (Sec. V.1) and the Full QDNN described in following section (Sec. V.2). The Basic QDNN is designed to mirror the CDNN model in architecture and complexity, serving as a baseline comparison without quantum-specific optimization. The Full QDNN is designed to fully take advantage of the quantum potential by carefully mitigating quantum-specific challenges and incorporating quantum-specific optimization schemes. The details of the architecture and construction of the CDNN and QDNNs, the method of extraction, and the corresponding results are presented in the following subsections.
@@ V.1 Basic QDNN
All the DNN models used for extraction in this paper have a three-dimensional input ($x_{B}$, $Q^{2}$, $t$), the kinematic values of a given kinematic bin, and a four-dimensional output ($\Re e\mathcal{H}$, $\Re e\mathcal{E}$, $\Re e\widetilde{\mathcal{H}}$, $DVCS$), the four CFFs. The methodology of the local CFF fitting used here follows a standard maximum-likelihood estimation approach except with the added benefit of leveraging neural networks optimization algorithms and flexibility. Such fits were first applied to local CFF extraction by Kumerički et al. [83]. For more detail on modern methodology and error propagation see [75]. The DNNs use a custom, physics-informed loss function that predicts the cross section $F$ using the outputted three CFFs and DVCS term with the BKM10 formalism (Sec. II) and computes the mean squared error of the predicted cross section against the true cross section.
Figure 1: Schematic of the CDNN used for local CFF regression. The network takes a three-dimensional kinematic input $(x_{B},Q^{2},t)$ and maps it through an input layer, eight fully connected hidden layers (64 neurons each) with ReLU activations, to a four-dimensional output $(\Re e\,\mathcal{H},\Re e\,\mathcal{E},\Re e\,\widetilde{\mathcal{H}},\mathrm{DVCS})$. Only the first and last hidden layers are shown explicitly; intermediate layers are omitted for clarity.
Figure 2: Schematic of the QDNN used for local CFF regression. A classical preprocessing layer maps $(x_{B},Q^{2},t)$ to $n=6$ qubits via angle embedding. The circuit applies $L=8$ Strongly Entangling Layers (SELs); only a subset is drawn for clarity. Pauli-$Z$ expectation values are measured on each qubit to form a 6-dimensional classical vector, which is passed to a small classical head to produce four outputs $(\Re e\,\mathcal{H},\Re e\,\mathcal{E},\Re e\,\widetilde{\mathcal{H}},\mathrm{DVCS})$.
The CDNN model used for the extraction of the CFFs is based closely on the CDNN regression feedforward architecture shown in Fig. 1. The input layer of the CDNN maps the three-dimensional input layer ($x_{B}$, $Q^{2}$, $t$) to 64 neurons. Then, there are 8 hidden layers, each with 64 neurons with rectified linear unit (ReLU) activation ($\mathrm{ReLU}(z)=\max(0,z)$). Finally, the output layer maps the 64 features to a four-dimensional output ($\Re e\mathcal{H}$, $\Re e\mathcal{E}$, $\Re e\widetilde{\mathcal{H}}$, $DVCS$). The QDNN is constructed based on the architecture of the CDNN and is similar to the architecture shown in Fig. 2. The input layer of the QDNN is a classical linear preprocessing layer, mapping the three-dimensional input to 6 qubits, which is equivalent to the 64 neurons used in the CDNN (Sec. III). The input data is then embedded into quantum state space using angle embedding, which encodes the features as rotation angles on the qubits. Next, there are 8 hidden Strongly Entangling Layers, consisting of rotations and entanglers. Finally, the expectation value of the Pauli-Z operator for each qubit is measured. In the classical postprocessing layers, the 6 quantum outputs are mapped to 64 neurons, followed by a Tanh activation, then another linear layer to the final four-dimensional output.
To enable a controlled benchmark, we keep the forward model, loss, training protocol, and input/output structure fixed, and choose architectures with comparable forward-pass FLOP estimates. The QDNN uses substantially fewer trainable parameters, so performance differences reflect model class and inductive bias rather than simply parameter count (Sec. V.1.1). Subsequently, we apply both models to extract Compton form factors (CFFs) from pseudodata and present a comprehensive error analysis of the results (Sec. V.1.2).
@@ V.1.1 Model Complexity
A fair comparison between the CDNN and QDNN requires ensuring that neither architecture gains an artificial advantage from simply being larger or more computationally powerful. We therefore quantify and compare two standard measures of model complexity—the number of trainable parameters and the approximate floating-point operation count (FLOPs) for a forward pass. These metrics characterize a model’s representational capacity and computational cost, respectively, and provide a baseline for assessing whether the CDNN and Basic QDNN operate at comparable levels of complexity despite belonging to fundamentally different paradigms.
Direct comparisons between classical and quantum models are inherently subtle. A QDNN may have far fewer tunable parameters, yet its entangling structure can encode highly non-linear transformations that would require a significantly larger classical network to approximate. Likewise, FLOPs are not strictly well-defined for quantum circuits, which evolve qubit states via unitary operations rather than classical arithmetic. Nevertheless, computing these metrics offers a useful reference point and ensures that neither architecture is advantaged by a disproportionate model size.
Trainable Parameters:
For a fully connected layer with input dimension $M$ and output dimension $N$, the number of trainable parameters is
$P=MN+N.$ (22)
The CDNN consists of an input layer ($3\!\to\!64$), eight hidden layers ($64\!\to\!64$), and an output layer ($64\!\to\!4$). The total number of trainable parameters is
$\begin{split}P_{\text{CDNN}}&=(3\cdot 64+64)+8(64\cdot 64+64)+(64\cdot 4+4)\\ &=33796.\end{split}$ (23)
The QDNN includes a preprocessing layer ($3\!\to\!6$), eight Strongly Entangling Layers (SELs), and postprocessing layers ($6\!\to\!64\!\to\!4$). Each SEL applies three parameterized rotations per qubit:
@@
The angle-embedding step contains no trainable parameters. The total parameter count is
$\begin{split}P_{\text{QDNN}}&=(3\cdot 6+6)+144+(6\cdot 64+64)+(64\cdot 4+4)\\ &=876.\end{split}$ (25)
As expected, the QDNN uses far fewer parameters due to its ability to encode a $2^{n}$-dimensional Hilbert space in $n$ qubits.
Floating-Point Operation Count:
For a classical dense layer, the forward-pass FLOP count is
@@
Including ReLU activations (64 FLOPs each), the total FLOPs are
$\begin{split}C_{\text{CDNN}}&=(2\cdot 3\cdot 64+64)+8(2\cdot 64\cdot 64+64)+(2\cdot 64\cdot 4)\\ &=67008.\end{split}$ (27)
Quantum layers act on the $2^{6}=64$-dimensional state vector. Approximating FLOPs for preprocessing, eight SELs, postprocessing layers, and a Tanh activation (64 FLOPs), we obtain
$\begin{split}C_{\text{QDNN}}&=(2\cdot 3\cdot 64)+8(2\cdot 64\cdot 64)\\ &\mathrel{\phantom{=}}+\bigl(2\cdot 64\cdot 64+64+2\cdot 64\cdot 4\bigr)\\ &=74688.\end{split}$ (28)
The resulting FLOP estimate is of the same order as the CDNN.
Table 3 summarizes the comparison. Despite their structural differences, the CDNN and QDNN exhibit similar effective complexity when assessed through parameter counts and approximate FLOPs, ensuring that our performance comparison is not biased by model size or computational load.
Figure 3: Distributions of the extracted $\Re e\mathcal{E}$ from noisy replicas of cross section pseudodata generated from kinematics $k=5.75$ GeV, $Q^{2}=2.22$ GeV2, $x_{B}=0.333$, t=0.16t=-0.16 GeV2 (Set 144). The histograms display improved accuracy and precision by the QDNN.
Figure 4: Predicted cross sections from the (a) CDNN and (b) QDNN fits of noisy replicas of cross section pseudodata (red points) generated from kinematics $k=8.521$ GeV, $Q^{2}=3.65$ GeV2, $x_{B}=0.367$, t=0.20459t=-0.20459 GeV2 (Set 26).
Figure 5: Predicted cross sections from the (a) CDNN and (b) QDNN fits of noisy replicas of cross section pseudodata (red points) generated from kinematics $k=5.75$ GeV, $Q^{2}=2.48$ GeV2, $x_{B}=0.399$, t=0.45t=-0.45 GeV2 (Set 165). In both cases there is a small improvement in proximity to the true with a clear narrowing of the error band for the QDNN.
@@ V.1.2 CFF Extraction Test
To quantify overall performance, we define a reduced $\chi^{2}$ metric, $M_{\chi^{2}}$, as the sum of the reduced $\chi^{2}$ values for each CFF across all 195 kinematic bins.
$M_{\chi^{2}}\equiv\sum_{F\in\{\Re e\mathcal{H},\Re e\mathcal{E},\Re e\widetilde{\mathcal{H}},\mathrm{DVCS}\}}\frac{1}{N_{\rm bins}}\sum_{b=1}^{N_{\rm bins}}\frac{\big(\bar{F}_{b}-F^{\rm true}_{b}\big)^{2}}{\sigma^{2}(F_{b})}.$ (29)
@@
A lower $M_{\chi^{2}}$ indicates a better fit, with a value of 4 corresponding to a perfect match. Averaging over all bins, we find $\overline{M}_{\chi^{2}}^{\text{CDNN}}\approx 78.97$ and $\overline{M}_{\chi^{2}}^{\text{QDNN}}\approx 28.18$, indicating that the QDNN outperforms the CDNN by approximately 64.3%, even without leveraging quantum-specific enhancements.
@@
The predicted cross sections using the predicted CFFs from the CDNN and QDNN are also analyzed. Using the framework described in Section II, we compute the predicted cross sections using the extracted CFFs and compare them to the true cross section values used to generate the pseudodata. In Figs. 4 and 5, we present results for two representative kinematic bins, plotting the mean predicted cross-section values as a function of the azimuthal angle $\phi$, along with their 68% confidence intervals. The true cross-section curve and fluctuated true cross-section points with simulated experimental error from the pseudodata sets are overlaid for reference. A key observation from these figures is that the improved precision of the QDNN extraction leads to narrower confidence intervals in the predicted cross sections across all three bins due to the reduced uncertainty in the extracted CFFs. In these examples, there is also a small, observable improvement in the accuracy of the predicted cross section by the QDNN.
@@
Understanding the sources of error is essential for robust model evaluation. A detailed analysis of the errors [75] in the CFF extractions is necessary to fully understand and quantify the differences between QDNN and CDNN approaches. We consider four distinct error metrics: algorithmic error, methodological error, precision, and accuracy. To compute the algorithmic error, we perform the extraction on 1000 identical replicas of each bin and calculate the resulting spread of each CFF. Thus, algorithmic error represents error resulting from the limitations and inherent uncertainties of the extraction algorithm itself, rather than simulated experimental error in the data. To compute the methodological error, we introduce a spread in the parameters used to generate the CFFs (Table 2) and use the generating function [Eq. (19)] to generate CFFs sampled in a domain that approximately corresponds to the spread in the original extracted CFFs. Then, the methodological error is computed as the spread of the resulting residuals from the new extractions. Methodological error represents the uncertainty arising from the choice of model parameters and assumptions used in the extraction process. By introducing a spread in the parameters used to generate the CFFs and propagating this spread through the generating function, we assess how variations in these underlying assumptions influence the extracted CFF values. This approach allows us to estimate the systematic impact of modeling choices on the final extraction results, distinguishing it from purely statistical or algorithmic uncertainties. Finally, precision and accuracy are computed using Eqs. (21) and (20), respectively. These metrics provide an evaluation of the performance of the DNNs on the original pseudodata set, with precision measuring the consistency of the extracted values across multiple replicas and accuracy reflecting the agreement between the extracted and true CFFs. For all four of these error metrics, a lower value indicates a lower error.
@@
Figure 6 presents a comparative analysis of algorithmic error, methodological error, precision, and accuracy for the DNN-based extraction of each CFF across all 195 kinematic bins. In the plot, the absolute mean deviation from the true value is shown for each contribution as a point, while the spread in that deviation is represented by the error bar for each point. A notable improvement in the three CFFs’ precision is observed with the QDNN, whereas the other error contributions remain largely comparable between the two models. The QDNN shows an improved extraction accuracy for $\Re e\mathcal{E}$ and $\Re e\widetilde{\mathcal{H}}$ while the CDNN performs better for $\Re e\mathcal{H}$ and $DVCS$. Methodological errors are fairly small in both cases, but with a narrower distribution for the QDNN, indicating a bit better robustness to variations in the model parameters used to generate the CFFs. The increase in precision from the QDNN is also evident in the predicted cross-section plots (Figs. 4 and 5) as visibly tighter confidence intervals. Overall, QDNN demonstrates improved total error performance compared to CDNN, underscoring its effectiveness in this application.